OpenStudy (anonymous):

What is the vertex of 3x^2+6x+7?

OpenStudy (anonymous):

Do you know the standard form of a parabola?

OpenStudy (anonymous):

a^2+b^2=c^2?

OpenStudy (anonymous):

No, that is a circle. For a parabola, it's y = a(x - h)^2 + k

OpenStudy (anonymous):

Where (h,k) is the vertex

OpenStudy (anonymous):

ahh okay thanks!

OpenStudy (anonymous):

Do you know how to put the equation into standard form?

OpenStudy (anonymous):

no sir

OpenStudy (anonymous):

Okay, do you know how to complete the square?

OpenStudy (anonymous):

so far all I have written down is a=3 b=6 and c=7 sorry, i'm really bad at math..

OpenStudy (anonymous):

I think you should brush up on your quadratics, but to complete the square, you have to first remove the coefficient in front of the leading term which is x^2. I would factor it out: \[3x^2+6x+7 = 3(x^2 + 2x + \frac{7}{3})\] complete the square: \[3[(x^2 + 2x +1) + \frac{7}{3} - 1] = 3[(x + 1)^2 + \frac{4}{3}]\]expand with the distributive property to get into the standard form: \[3[(x + 1)^2 + \frac{4}{3}] = 3(x+1)^2 + 4\] and now you can identify your vertex (h,k) quite easily: (-1,4)